# Probing Picohertz Gravitational Waves with Pulsars

### v1

@@ V Results
We then search for CW signals in the real data, running up to $3,000,000$ samples in MCMC and discard the first $20\%$ as burn-in. For the search we use log uniform prior on $h_{0}$. No CW signals are found in either data set. Fig. 4 and Fig. 5 show the posterior distributions of the relevant parameters for $\dot{P}_{b}$ search and $\ddot{P}$ search respectively. We recover the priors from the analysis except for a constrained strain amplitude. We report $BF=0.6913\pm 0.0058$ for the $\dot{P}_{b}$ search and $BF=0.4627\pm 0.0042$ for the $\ddot{P}$ search. Based on the null distribution and the posterior distributions, we do not detect a signal.
Figure 4: $\dot{P}_{b}$ search on current data with uniform prior of $f_{\text{GW}}$. Upper limit on $h_{0}$ can be extracted from the posterior distribution. The priors (blue dashed line) are also shown. The contours are set at $50\%,80\%,95\%$. The following figures follow the same convention.
@@ V.1 Upper limits
With our new analytic model, Bayesian framework, and the updated $\dot{P}_{b}$ data, we improve the $h_{0}$ upper limits by 9.68 times averaging over frequency, approximately an order of magnitude over previous results [27]. For the current data, the $\ddot{P}$ analysis achieves tighter strain limits than $\dot{P}_{b}$ for $f_{\text{GW}}\gtrsim 1.5~\text{pHz}$. Nevertheless, a combined analysis across multiple timing parameters remains valuable: when the sensitivity reaches levels comparable to expected CW signals, joint analysis of $\dot{P}_{b}$, $\ddot{P}$, and potentially other parameters could break degeneracies in source properties including CW frequency and sky locations, providing stronger constraints on strain amplitude.
On top of the fully marginalized upper limit, we also calculate the sky location dependent upper limit, presented in Fig. 7. This is done by fixing $l,b$ in the Bayesian analysis for the corresponding sky location. In case of no detection, we expect the sensitivity distribution to correlate with the pulsar distribution by virtue of the antenna pattern function defined in Eq. 2. Our result conforms to this expectation. In the calculation presented here, we use a reference GW frequency of $100~\text{pHz}$. We note that the detection sensitivity can be different by up to an order of magnitude in different regions of sky. We achieve as best as $h_{0}\simeq 6.7\times 10^{-13}$ in $95~\%$ upper limit for particular sky locations, one seventh of the marginalized result over all sky at the same frequency.
@@ V.2 Scaling laws
Figure 6: SMBHB inverse residence time as a function of orbital frequencies, for different total masses and mass ratios. The turnover at $\sim\text{nHz}$ corresponds to the turnover in the GWB spectrum in Fig. 2. LHS: inverse residence time as a function of total SMBHB mass, where mass ratio =1=1 for all SMBHBs; RHS: inverse residence time as a function of SMBHB mass ratio, where $M_{\text{BH, 12}}=10^{9}~\mathrm{{M}}_{{\odot}}$.
Figure 7: $95\%$ upper limit on $h_{0}$ as a function of sky location, at a reference frequency of $100\text{pHz}$. Left: $\dot{P}_{b}$ data. The highest $h_{0}$ upper limit is $5.1\times 10^{-10}$ and the lowest upper limit is $1.1\times 10^{-10}$; Right: $\ddot{P}$ data. The highest $h_{0}$ upper limit is $5.4\times 10^{-12}$ and the lowest upper limit is $6.7\times 10^{-13}$. The red diamonds denote the locations of the pulsars in the sky. The sensitivity roughly follows the antenna beam pattern.
Figure 8: The number of SMBHBs per 0.2 log frequency (Hz) as a function of mass range. In pHz regime, SMBHBs across the mass range are likely to contribute to CW signal. The SMBHB population is computed by modifying oddAGN [40].
@@ V.3 SMBHB synthesis model
The integrated GWB spectrum predicted from our SMBHB population model deviates significantly from a simple power-law extrapolation in the sub-nHz regime. In particular, a turnover in the spectrum appears around $60~\text{pHz}$, corresponding to the strongest GWB signal across the entire frequency range. A simple power-law assumes that GW emission dominates SMBHB orbital evolution, while our model incorporates environmental effects such as stellar hardening and dynamical friction. At pHz frequencies, these environmental processes dominate, shortening the binary residence timescale ($(df_{\mathrm{orb}}/dt)^{-1}$) compared to purely GW-driven evolution (Fig. 6), where $f_{\mathrm{orb}}$ is the binary orbital frequency. The shorter residence time leads to a suppression of GWB power at the lowest frequencies, manifested as a reduced characteristic strain in Fig. 2.
@@
We extract the loudest individual CW sources in each frequency bin. We find no strong scaling between the strain amplitude of the loudest sources and frequency. To extract these sources, we mask all the individual bins in the parameter space with at least one SMBHB, and take the largest strain amplitude for each frequency bin. The loudest sources typically exhibit strain amplitudes around $1\sim 4\times 10^{-16}$. The majority of individual sources lie several orders of magnitude below this level.
@@ V.4 Future Detection Prospects
To forecast detection prospects, we conduct simulations to determine the scaling relations between detection sensitivity and (1) data size and (2) data uncertainty levels. To generate the simulated pulsar data, we draw from the extrapolated distribution of pulsar location based on [41]. We verify that, when probing strain amplitudes above potential CW sources, uncertainties in pulsar distances alone do not significantly impact the upper limits on $h_{0}$. Each simulation assumes a null CW signal. We find the detection sensitivity scales as $\propto\sigma_{a}/\sqrt{N}$ when $N$ is asymptotically large (above $\sim 100$), where $\sigma_{a}$ denotes the uncertainty in data used for inference and $N$ is the number of pulsars (Fig. 9). When $N$ is below 50, boundary effects kick in and the scaling deviate from $\propto 1/\sqrt{N}$. In fact, we find the improvement of $h_{0}$ upper limit with respect to $N$ to be more efficient when $N$ is below 50.
Figure 9: Scaling of detection sensitivity with data size N, calculated at a reference frequency of $10~\text{pHz}$. The fitted lines have slopes of roughly -0.5, which agrees with the expected scaling law with N.
@@
Upcoming experiments will dramatically expand the PTA by increasing the number of precisely timed MSPs to the order of several hundred, representing more than an order-of-magnitude increase in size relative to current data [41]. Distance measurements will also improve significantly with SKA astrometry [42], reducing uncertainties to levels that satisfy our distance-precision criterion for essentially all pulsars. According to [41], SKA1 is expected to produce precise measurements for 675 pulsars, with an additional 120 achieved by SKA2.
@@
Figure 10: Forecast sky-location dependent strain upper limits with SKA. The sky locations of the pulsars are drawn from a distribution extrapolated based on current data. The highest $h_{0}$ upper limit is $6.8\times 10^{-13}$ and the lowest upper limit is $1.4\times 10^{-13}$.
@@
For a conservative forecast, we assume each pulsar retains the typical uncertainty level seen in the current data. We further assume that $80\%$ of the $\sim 700$ precisely timed pulsars are in binary systems [43], each with a measurable $\dot{P}_{b}$, and that all have measurable $\ddot{P}$. In addition, we take the observation time for all the new pulsars to be 10 years, which is below the longest observation time in the current data. Under these assumptions, we compare the forecast sensitivities derived from $\dot{P}_{b}$ and $\ddot{P}$ analyses with the strain amplitudes of the loudest individual sources from our SMBHB population model. As shown in Fig. 2, SKA-based data will enable an additional order-of-magnitude improvement over current sensitivities, solely due to the increase in data size. Our simulations also show that with more pulsars, the detection sensitivity will become more isotropic (Fig. 10). SKA data is expected to reduce the variance between the most and least sensitive sky locations by $\sim 40\%$ for $\ddot{P}$ analysis. However, the strain amplitudes of individual SMBHB CW sources drawn from our GWB synthesis model are still $\sim 3$ orders of magnitude below this forecast sensitivity.
@@ V.5 Injection of a CW
Figure 11: The scaling between Bayes factors and number of pulsars in the simulated data. The results being plotted are for $\ddot{P}$.
Figure 12: Two strategies to achieve a detection: (a) Improve the precision of $\ddot{P}$ measurement for all the pulsars; (b) Improve the precision of $\ddot{P}$ measurement for a selected few pulsars. Simulations show that an improvement of 2.5 (3) orders of magnitude is necessary for a detection in the former (latter) case in data uncertainty.
In our simulations, we choose NGC 6240 as the host galaxy of the SMBHB. NGC 6240 is a nearby ultraluminous infrared galaxy and prominent candidate for dual AGN, which in turn strongly indicates the presence of a SMBHB [44]. We assume the CW from the SMBHB has the strongest strain amplitude predicted by our population model and a redshifted frequency of $300~\text{pHz}$.
@@
To forecast detection prospects with improving data uncertainty, we carry out simulations under two scenarios, as shown in Fig. 12. The first is when the parameter measurements for all the pulsars improve at the same rate, plotted as a black solid line, while the second is considering a scenario where 10 pulsars have improving data uncertainty and the remaining pulsars have their precision fixed at the typical uncertainty in current data. Our results suggest a uniform improvement of two orders of magnitude on the existing sensitivity ($\delta\ddot{P}/P\sim\times 10^{-30}s^{-2}$) is necessary, or, alternatively a three order of magnitude improvement for 10 out of the 795 pulsars. So there is a clear trade-off between precision required and number of pulsars that can achieve this precision—the larger the number of pulsars achieving the precision, the lower the precision requirement. Though it may seem as if a uniform improvement of two orders of magnitude would be easier to achieve, it may actually prove the case that an improvement on only 10 pulsars is more realistic. Measuring $\ddot{P}$ requires extremely stable pulsars and long timing baselines, hence it is likely that even with SKA data, only a select few pulsars will be able to provide precise measurements of their $\ddot{P}$.
@@ VI Discussion and Conclusions
As difficult as the required improvement in timing parameter measurement for a future detection may seem, multiple factors work in our favor. Current studies indicate the timing precision of SKA could be up to 2 orders of magnitude better than today [45]. On top of that, it has been empirically shown that a longer observation time reduces the standard deviations in timing parameters [46], and simulations suggest a scaling of $\sim T_{\mathrm{obs}}^{-1/2}$ [47]. Hence it is possible that some potential CW sources may fall within our sensitivity range with SKA data.
@@
Currently, the dominant sources of uncertainty in both $\dot{P}_{b}$ and $\ddot{P}$ are due to the measured PTA timing parameters, which are expected to improve with longer time baselines and next-generation telescopes such as SKA. As timing precision improves, however, red noise and uncertainties in galactic acceleration will no longer be sub-dominant. These effects present their own challenges and motivate the exploration of higher-order timing parameters with better-controlled uncertainties. We also acknowledge that only a very small fraction of pulsars will have precise measurements of higher-order timing parameters, due to the stringent requirement of pulsar stability and long timing baseline. However, we have shown in Sec. V.5 that a small number of precise measurements could potentially facilitate a successful detection.
@@
Targeted search may also help improve detection prospects. In our framework, targeted search is achieved by adopting informative priors on CW parameters. These priors are often informed by multi-messenger astrophysical observations. For example, recent work has shown that dual AGNs can serve as tracers of SMBHBs, where the sky location of candidate dual AGNs can be determined through optical, radio, and X-ray observations, while redshifts are measured spectroscopically [29]. Moreover, the orientation of SMBHB candidates can be constrained through observations of jet alignment [48], broad-line region kinematics [49], variability and polarization of quasar light [50], and host galaxy morphology [51].
@@
The framework presented here can also be used to search for alternative deterministic GW signals by adapting the CW model parameters. This opens opportunities to search for a variety of exotic physics, including ultralight dark matter and dark matter substructure [52, 53].
@@
The exact effect of gas hardening on binary formation and evolution is still an open research question. For our GWB synthesis model, the energy loss of the binary by gas hardening is likely to be comparable with GW emission and stellar hardening at a transitional binary separation of $\sim 1$ pc [54, 55]. Below this separation, the energy loss is dominated by GW emission; above this separation, energy loss is dominated by stellar hardening. Inclusion of gas hardening in the model may increase the turnover frequency in the GWB spectrum, since it will further reduce the residence time of a binary and thus GWs emitted at a certain frequency. Towards the very beginning of binary formation, gas hardening is expected to be subdominant to stellar hardening and dynamical friction, so the low-frequency asymptotic behavior of our GWB spectrum should remain unaltered.

### v2

@@ V Results
We then search for CW signals in the real data, running up to $3,000,000$ samples in MCMC and discarding the first $20\%$ as burn-in. For the search we use log uniform prior on $h_{0}$. No CW signals are found in either data set. Fig. 4 and Fig. 5 show the posterior distributions of the relevant parameters for $\dot{P}_{b}$ search and $\ddot{P}$ search respectively. We recover the priors from the analysis except for a constrained strain amplitude. We report $BF=0.6913\pm 0.0058$ for the $\dot{P}_{b}$ search and $BF=0.4728\pm 0.0042$ for the $\ddot{P}$ search. Based on the null distribution and the posterior distributions, we do not detect a signal.
Figure 4: $\dot{P}_{b}$ search on current data with uniform prior of $f_{\text{GW}}$. Upper limit on $h_{0}$ can be extracted from the posterior distribution. The priors (blue dashed line) are also shown. The contours are set at $50\%,80\%,95\%$. All the angles and phase are in radians, and $f_{\text{GW}}$ is in $\log_{10}\text{Hz}$. The following figures follow the same convention.
@@ V.1 Upper limits
With our new analytic model, Bayesian framework, and the updated $\dot{P}_{b}$ data, we improve the $h_{0}$ upper limits by 9.53 times at a reference frequency of $f_{\text{GW}}=100~\text{pHz}$, an order of magnitude over previous results [DeRocco2023]. For the current data, the $\ddot{P}$ analysis achieves tighter strain limits than $\dot{P}_{b}$ for $f_{\text{GW}}\gtrsim 1.5~\text{pHz}$. Nevertheless, a combined analysis across multiple timing parameters remains valuable: when the sensitivity reaches levels comparable to expected CW signals, joint analysis of $\dot{P}_{b}$, $\ddot{P}$, and potentially other parameters could break degeneracies in source properties including CW frequency and sky locations, providing stronger constraints on strain amplitude.
On top of the fully marginalized upper limit, we also calculate the sky location dependent upper limit, presented in Fig. 7. This is done by fixing $\alpha,\,\delta$ in the Bayesian analysis for the corresponding sky location. In case of no detection, we expect the sensitivity distribution to correlate with the pulsar distribution by virtue of the antenna pattern function defined in Eq. 2. Our result conforms to this expectation. In the calculation presented here, we use a reference GW frequency of $100~\text{pHz}$. We note that the detection sensitivity can be different by up to an order of magnitude in different regions of sky. The upper limit in strain is as low as $h_{0}\simeq 6.7\times 10^{-13}$ for particular sky locations, one seventh of the marginalized result over all sky at the same frequency.
@@ V.2 Scaling laws
Figure 6: SMBHB residence time $T_{\text{res}}=f_{\text{orb}}/\dot{f}_{\text{orb}}$ as a function of orbital frequencies, for different total masses and mass ratios. The turnover at $\sim\text{nHz}$ corresponds to the turnover in the GWB spectrum in Fig. 2. LHS: residence time as a function of total SMBHB mass, where mass ratio =1=1 for all SMBHBs; RHS: residence time as a function of SMBHB mass ratio, where $M_{\text{SMBHB}}=10^{9}~\mathrm{{M}}_{{\odot}}$.
Figure 7: $95\%$ upper limit on $h_{0}$ as a function of sky location, at a reference frequency of $100~\text{pHz}$. Left: $\dot{P}_{b}$ data. The highest $h_{0}$ upper limit is $5.1\times 10^{-10}$ and the lowest upper limit is $1.1\times 10^{-10}$; Right: $\ddot{P}$ data. The highest $h_{0}$ upper limit is $5.4\times 10^{-12}$ and the lowest upper limit is $6.7\times 10^{-13}$. The red diamonds denote the locations of the pulsars in the sky. The sensitivity roughly follows the antenna beam pattern.
Figure 8: The number of SMBHBs per 0.2 log frequency (Hz) as a function of mass range. In pHz regime, SMBHBs across the mass range are likely to contribute to CWs. The SMBHB population is computed by modifying oddAGN [oddAGN]. There is a turn-over in $N_{\text{SMBHB}}$ at pHz frequencies for all mass ranges considered.
@@ V.3 SMBHB synthesis model
The integrated GWB spectrum predicted from our SMBHB population model deviates significantly from a simple power-law extrapolation in the sub-nHz regime. In particular, a turnover in the spectrum appears around $60~\text{pHz}$, corresponding to the strongest GWB signal across the entire frequency range. A simple power-law assumes that GW emission dominates SMBHB orbital evolution, while our model incorporates environmental effects such as stellar hardening and dynamical friction. At pHz frequencies, these environmental processes dominate, shortening the binary residence timescale ($f_{\mathrm{orb}}/\dot{f}_{\mathrm{orb}}$) compared to purely GW-driven evolution (Fig. 6), where $f_{\mathrm{orb}}$ is the binary orbital frequency. The shorter residence time leads to a suppression of GWB power at the lowest frequencies, manifested as a reduced characteristic strain in Fig. 2.
@@
We extract the loudest individual CW sources in each frequency bin. We find no strong scaling between the strain amplitude of the loudest sources and frequency. To extract these sources, we find the largest strain amplitude from all the bins that contains a SMBHB, for each frequency. The loudest sources typically exhibit strain amplitudes around $1\sim 4\times 10^{-16}$. The majority of individual sources lie several orders of magnitude below this level.
@@ V.4 Future Detection Prospects
To forecast detection prospects, we conduct simulations to determine the scaling relations between detection sensitivity and (1) data size and (2) data uncertainty levels. To generate the simulated pulsar data, we draw from the extrapolated sky location distribution of known pulsars inspired by method described in [Xin2021]. Both conservative and more optimistic estimates for SKA1 and 2 are included here. The SKA LOW survey will concentrate on regions within $\pm 5^{\circ}$ from the galactic plane, while the SKA MID survey will concentrate on regions within $\pm 10^{\circ}$ of the galactic plane [Braun:2015zta]. Upcoming experiments will dramatically expand the PTA by increasing the number of precisely timed MSPs to the order of several hundred, representing more than an order-of-magnitude increase in size relative to current data [Xin2021]. Distance measurements will also improve significantly with SKA astrometry [Smits:2011], reducing uncertainties to levels that satisfy our distance-precision criterion for essentially all pulsars. According to [Xin2021], SKA1 is expected to produce precise measurements for 675 pulsars, with an additional 120 achieved by SKA2. Because SKA1 consists of southern hemisphere telescopes, we apply a cut-off at $30^{\circ}$ in the declination angle of the pulsars. We then draw pulsar locations from our empirical sky location distribution extracted from all the pulsars in IPTA data [Antoniadis2022].
We verify that, when probing strain amplitudes above potential CW sources, uncertainties in pulsar distances alone do not significantly impact the upper limits on $h_{0}$. Each simulation assumes a null CW signal. We find the detection sensitivity scales as $\propto\sigma_{p}/\sqrt{N}$ when $N$ is asymptotically large (above $\sim 100$), where $\sigma_{p}$ denotes the uncertainty in data used for inference and $N$ is the number of pulsars (Fig. 9). When $N$ is below 50, boundary effects kick in and the scaling deviates from $\propto 1/\sqrt{N}$—the exponent on $N$ in the power law scaling is slightly smaller than 1/2-1/2.
Figure 9: Scaling of detection sensitivity measured by detectable CW strain amplitude as a function of data size N, calculated at a reference frequency of $10~\text{pHz}$ as a conservative estimate, for both $\dot{P}_{b}$ and $\ddot{P}$ analyses. For higher frequencies, the sensitivity improves. For $\dot{P}_{B}$ analysis, $\sigma_{p}$ denotes the uncertainty in $\Delta\dot{P}_{b}/P_{b}$, which is the observed quantity that should be equal to the CW dependent $a_{\text{GW}}$; for $\ddot{P}$ analysis, $\sigma_{p}$ denotes the uncertainty in $\ddot{P}/P$, which is the observed quantity that should be equal to the CW dependent $j_{\text{GW}}$. The fitted lines have slopes of roughly -0.5, which agrees with the expected scaling law with N. For any larger frequencies, the method is sensitive to lower strain amplitude: for $\dot{P}_{b}$ analysis, $h_{0}\propto f_{\text{GW}}^{-1}$; for $\ddot{P}$ analysis, $h_{0}\propto f_{\text{GW}}^{-2}$
@@
Figure 10: Forecast sky-location dependent strain upper limits in SKA era at a reference frequency $f_{\text{GW}}=100~\text{pHz}$. On top of the pulsars already timed by EPTA, NANOGrav and PPTA, the sky locations of the SKA pulsars are drawn from a distribution extrapolated based on currently known pulsars. The highest $h_{0}$ upper limit is $7.8\times 10^{-13}$ and the lowest upper limit is $1.3\times 10^{-13}$.
@@
For a conservative forecast, we assume each pulsar retains the typical uncertainty level seen in the current data. We further assume that $80\%$ of the $\sim 800$ precisely timed pulsars are in binary systems [Lorimer2008], each with a measurable $\dot{P}_{b}$, and that all have measurable $\ddot{P}$. In addition, we take the observation time for all the new pulsars to be 10 years, which is below the longest observation time in the current data. Under these assumptions, we compare the forecast sensitivities derived from $\dot{P}_{b}$ and $\ddot{P}$ analyses with the strain amplitudes of the loudest individual sources from our SMBHB population model. As shown in Fig. 2, SKA-based data will enable an additional order-of-magnitude improvement over current sensitivities, solely due to the increase in data size. Our simulations also show that with more pulsars, the detection sensitivity will become more isotropic (Fig. 10). SKA data is expected to reduce the variance between the most and least sensitive sky locations by $\sim 40\%$ for $\ddot{P}$ analysis. However, the strain amplitudes of individual SMBHB CW sources drawn from our GWB synthesis model are still $\sim 3$ orders of magnitude below this forecast sensitivity.
@@ V.5 Injection of a CW
Figure 11: The scaling between Bayes factors favoring the existence of the injected CW signal over null and number of pulsars N in the simulated data. The results being plotted are for $\ddot{P}$ analysis, which is more promising for a detection. The detection threshold corresponds to the $95\%$ cut-off in the null BF distribution. $\sigma_{p}$ is the typical uncertainty in the measured $\ddot{P}/P$ for a pulsar in the dataset.
Figure 12: Two strategies to achieve a detection: (a) Improve the precision of $\ddot{P}$ measurement for all the pulsars; (b) Improve the precision of $\ddot{P}$ measurement for a selected few pulsars. Simulations show that an improvement of 2.5 (3) orders of magnitude is necessary for a detection in the former (latter) case in data uncertainty. The detection threshold is set by the $95\%$ cut-off in the BF null distribution.
In our simulations, we choose NGC 6240 as the host galaxy of the SMBHB. NGC 6240 is a nearby ultraluminous infrared galaxy and prominent candidate for dual AGN, which in turn strongly indicates the presence of a SMBHB [sobolenko2022ngc6240]. We assume the CW from the SMBHB has the strongest strain amplitude predicted by our population model and a redshifted frequency of $300~\text{pHz}$.
@@
To forecast detection prospects with improving data uncertainty, we carry out simulations under two scenarios, as shown in Fig. 12. The first is when the parameter measurements for all the pulsars improve at the same rate, plotted as a black solid line, while the second is considering a scenario where 10 pulsars have improving data uncertainty and the remaining pulsars have their precision fixed at the typical uncertainty in current data. Our results suggest a uniform improvement of two orders of magnitude on the existing sensitivity ($\delta(\ddot{P}/P)\sim\times 10^{-30}s^{-2}$) is necessary, or, alternatively a three order of magnitude improvement for 10 out of the 795 pulsars. So there is a clear trade-off between precision required and number of pulsars that can achieve this precision—the larger the number of pulsars achieving the precision, the lower the precision requirement. Though it may seem as if a uniform improvement of two orders of magnitude would be easier to achieve, it may actually prove the case that an improvement on only 10 pulsars is more realistic. Measuring $\ddot{P}$ requires extremely stable pulsars and long timing baselines, hence it is likely that even with SKA data, only a select few pulsars will be able to provide precise measurements of their $\ddot{P}$.
@@ VI Discussion and Conclusions
As difficult as the required improvement in timing parameter measurement for a future detection may seem, multiple factors work in our favor. Current studies indicate the timing precision of SKA could be up to 2 orders of magnitude better than today [liu2014pulsar]. On top of that, it has been empirically shown that a longer observation time reduces the standard deviations in timing parameters [andrews2020nanograv], and simulations suggest a scaling of $\sim T_{\mathrm{obs}}^{-1/2}$ [liu2019correlated]. The typical timing precision of current operational PTAs is $\sim 1~\mu s$, and the current uncertainty of the period normalized $\dot{P}_{b}/P_{b}$, $\ddot{P}/P$ sits at $10^{-20}s^{-1}$, $10^{-30}s^{-2}$. Optimistically the uncertainties in the timing parameters scale linearly with the timing precision. With 2 orders or magnitude improvement in the timing precision, we would achieve $\sigma(\dot{P}_{b}/P_{b})\sim 10^{-22}s^{-1}$  /  $\sigma(\ddot{P}/P)\sim 10^{-32}s^{-2}$. The observation times of current PTAs range from a few years to two decades. Longer timespan will only improve the uncertainty in timing parameters, though the improvement due to longer timespan will be incremental compared to the improvement due to better timing precision by the scaling relations. Referring to Fig 12, it is likely that some potential CW sources may fall within our sensitivity range with SKA data.
@@
However, there are a few caveats. Currently, the dominant sources of uncertainty in both $\dot{P}_{b}$ and $\ddot{P}$ are due to the measured PTA timing parameters, which are expected to improve with longer time baselines and next-generation telescopes such as SKA. As timing precision improves, however, red noise and uncertainties in galactic acceleration will not be negligible. These effects present their own challenges and motivate the exploration of higher-order timing parameters with better-controlled uncertainties. More details of recent developments can be found for red noise mitigation in [Goncharov_2020] and for Galactic acceleration modeling in [ibata2023chartinggalacticaccelerationfield]. In particular, since red noise is covariant with a GWB signal, it will be interesting to include a GWB when extracting timing parameters, which has not been done. We also acknowledge that only a very small fraction of pulsars will have precise measurements of higher-order timing parameters, due to the stringent requirement of pulsar stability and long timing baseline. However, we have shown in Sec. V.5 that a small number of precise measurements could potentially facilitate a successful detection.
@@
Targeted searches may also help improve detection prospects [Agarwal:2025cag]. In our framework, targeted searches are achieved by adopting informative priors on CW parameters. These priors are often informed by multi-messenger astrophysical observations. For example, recent work has shown that dual AGNs can serve as tracers of SMBHBs, where the sky location of candidate dual AGNs can be determined through optical, radio, and X-ray observations, while redshifts are measured spectroscopically [Casey-Clyde2022]. Moreover, the orientation of SMBHB candidates can be constrained through observations of jet alignment [Dotti2010], broad-line region kinematics [Afanasiev2019], variability and polarization of quasar light [Hutsemekers2005], and host galaxy morphology [Zakamska2006].
@@
The framework presented here can also be used to search for alternative deterministic GW signals by adapting the CW model parameters. This opens opportunities to search for a variety of exotic physics, including ultralight dark matter and dark matter substructure [Brito2015, NANOGrav2023NewPhysics].
@@
The exact effect of gas hardening on binary formation and evolution is still an open research question. For our GWB synthesis model, the energy loss of the binary by gas hardening is likely to be comparable with GW emission and stellar hardening at a transitional binary separation of $\sim 1$ pc [Koss2023UGC4211, Izquierdo‐Villalba2021]. Below this separation, the energy loss is dominated by GW emission; above this separation, energy loss is dominated by stellar hardening. Inclusion of gas hardening in the model may increase the turnover frequency in the GWB spectrum, since it will further reduce the residence time of a binary and thus GWs emitted at a certain frequency. Towards the very beginning of binary formation, gas hardening is expected to be subdominant to stellar hardening and dynamical friction, so the low-frequency asymptotic behavior of our GWB spectrum should remain unaltered.